# Algebraic depletion interactions in two-temperature mixtures

### v1

@@ Acknowledgements.
The authors acknowledge fruitful discussions with Denis Dumont and Cesare Nardini and the support by F.R.S.-FNRS under the research Grant (PDR ”Active Matter in Harmonic Trap”) No. T.0251.20.
@@ Appendix A Details of the numerical simulations
The timestep was adjusted to ensure the stability and reproducibility of the simulations. The interaction between all particles were described with the pair style command and a purely repulsive harmonic-cut potential, $E=\epsilon(r_{c}-r)^{2}$ for $r<r_{c}$ and 0 everywhere else. The particle size was set equal to $r_{c}=1$ and various potential strengths were used and are indicated in the manuscript. This repulsive potential ensures a optimal stability of the simulations. Finally, to improve the accuracy of the data, the $g(r)$ were averaged on a large number of simulations (1000 to 10 000range100010000100010\,000start_ARG 1000 end_ARG to start_ARG 10 000 end_ARG). This was mandatory to have reliable results for the algebraic decay of the correlations.
@@ C.2.2 Solution
Finally, using these relations in Eqs. (52, 53), we obtain
@@ D.3.2 Order $\epsilon^{2}$
Using the decomposition (87), we have to determine the integral of $w_{123}$, $w_{231}$ and $w_{312}$ over $\bm{x}_{3}$:
$\int w_{123}\text{d}\bm{x}_{3}$ $=\int\omega_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (110)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}_{123}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{2}-\bm{x}_{1})+i\bm{k}^{\prime}\cdot(\bm{x}_{3}-\bm{x}_{1})}\text{d}\bm{x}_{3}$ (111)
$=\int\frac{\text{d}\bm{k}}{(2\pi)^{d}}\tilde{\omega}_{123}(\bm{k},0)e^{i\bm{k}\cdot(\bm{x}_{2}-\bm{x}_{1})}$ (112)
@@
$\int w_{231}\text{d}\bm{x}_{3}$ $=\int\omega_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (114)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}_{231}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{3}-\bm{x}_{2})+i\bm{k}^{\prime}\cdot(\bm{x}_{1}-\bm{x}_{2})}\text{d}\bm{x}_{3}$ (115)
$=\int\frac{\text{d}\bm{k}^{\prime}}{(2\pi)^{d}}\tilde{\omega}_{231}(0,\bm{k}^{\prime})e^{i\bm{k}^{\prime}\cdot(\bm{x}_{1}-\bm{x}_{2})}$ (116)
@@
$\int w_{312}\text{d}\bm{x}_{3}$ $=\int\omega_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (118)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}_{312}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{1}-\bm{x}_{3})+i\bm{k}^{\prime}\cdot(\bm{x}_{2}-\bm{x}_{3})}\text{d}\bm{x}_{3}$ (119)
$=\int\frac{\text{d}\bm{k}}{(2\pi)^{d}}\tilde{\omega}_{312}(\bm{k},-\bm{k})e^{i\bm{k}\cdot(\bm{x}_{1}-\bm{x}_{2})}$ (120)
@@
$\hat{g}_{2,2}=-\int\phi^{(2)}\text{d}\bm{x}_{3}=-\tau_{312}\int w_{312}\text{d}\bm{x}_{3}=\frac{\tau_{312}}{2T_{12}}v_{31}*v_{32}(\bm{x}_{1}-\bm{x}_{2}).$ (123)
@@ D.3.3 Order $\epsilon^{3}$
Using the decomposition (87), we have to determine the integral of the six terms $v_{13}w_{123}^{(2)}$, etc. over $\bm{x}_{3}$. They are (we omit the superscript “$(2)$” and use $\bm{x}=\bm{x}_{1}-\bm{x}_{2}$)
$\int v_{13}w_{123}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (127)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}_{123}(\bm{k},\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (128)
@@
$\int v_{23}w_{123}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (129)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k})\tilde{\omega}_{123}(\bm{k}^{\prime},\bm{k})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (130)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}_{123}(\bm{k}-\bm{k}^{\prime},\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (131)
@@
$\int v_{13}w_{231}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (132)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k})\tilde{\omega}_{231}(\bm{k},\bm{k}^{\prime})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (133)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}_{231}(\bm{k}^{\prime},\bm{k}-\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (134)
@@
$\int v_{23}w_{231}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (135)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}_{231}(\bm{k}^{\prime},\bm{k})e^{i\bm{k}\cdot\bm{x}}.$ (136)
@@
$\int v_{13}w_{312}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (137)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k})\tilde{\omega}_{312}(\bm{k}^{\prime},-\bm{k}-\bm{k}^{\prime})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (138)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}_{312}(-\bm{k}-\bm{k}^{\prime},\bm{k})e^{i\bm{k}\cdot\bm{x}}.$ (139)
@@
$\int v_{23}w_{312}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (140)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k})\tilde{\omega}_{312}(\bm{k}^{\prime},-\bm{k}-\bm{k}^{\prime})e^{i\bm{k}^{\prime}\cdot\bm{x}}$ (141)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}_{312}(\bm{k},-\bm{k}-\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (142)
@@ D.4 Scaling relations and numerical integration
D.4.1 Scaling relation
All the terms that involve the many-body potential $\phi$ at order $\epsilon^{3}$ consist of a double integral in Fourier space, three interaction potentials and a scale-invariant combination of the wavevectors (which scales as $k^{0}$). While the prefactor is difficult to obtain analytically, this form suggests an algebraic decay as $r^{-2d}$, where $d$ is the spatial dimension, and a prefactor scaling as $\tilde{V}(0)^{3}$, where $\tilde{V}(0)=\int V(\bm{x})\text{d}\bm{x}$.

### v2

@@ Acknowledgements.
The authors acknowledge fruitful discussions with Denis Dumont and Cesare Nardini and the support by F.R.S.-FNRS under the research Grant (PDR ”Active Matter in Harmonic Trap”) No. T.0251.20. This work was supported by the Belgian National Fund for Scientific Research (FRS-FNRS) within the Consortium des équipements de Calcul Intensif – CECI, under Grant 2.5020.11t. Quentin Thomas is FRIA grant holder from the Belgian National Fund for Scientific Research (FRS-FNRS).
@@ Appendix A Details of the numerical simulations
The timestep was adjusted to ensure the stability and reproducibility of the simulations. The interaction between all particles were described with the pair style command and the potentials given in the manuscript. Finally, to improve the accuracy of the data, the correlation $\hat{g}(r)$ has been averaged on a large number of simulations (1000 to 10 000range100010000100010\,000start_ARG 1000 end_ARG to start_ARG 10 000 end_ARG). This was mandatory to have reliable results for the algebraic decay of the correlations.
Four potentials were used for the pair interactions between particules:
  • Weeks-Chandler-Andersen (WCA) and Lennard-Jones (LJ),
    $U(r)=4u_{0}\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right]\quad r<r_{c}$ (20)
    To obtain the purely repulsive WCA potential, the value of $r_{c}$ is fixed at $\sigma 2^{1/6}$. For the classical LJ potential, we consider $r_{c}\gg\sigma$ to get the attractive part of the potential. It is referred as LJ/cut in LAMMPS.
  • Harmonic,
    $U(r)=\kappa(r_{c}-r)^{2}\quad r<r_{c}$ (21)
    In the simulations, we choose $r_{c}=1$. This potential is referred as harmonic/cut in LAMMPS.
  • Gaussian (gauss/cut in LAMMPS),
    $U(r)=\frac{H}{\sigma\sqrt{2\pi}}\exp\left[-\frac{(r-r_{m})^{2}}{2\sigma^{2}}\right]$ (22)
    The peak position, $r_{m}$, was fixed at zero. The global cutoff was chosen to obtain a convergence of the $g(r)$ correlation function computed from the simulation results.
In the manuscript, we use $\epsilon$ to describe the magnitude of the interaction potential. For WCA/LJ, harmonic and Gaussian, this parameter is defined as $\epsilon=4u_{0},\kappa,H/\sigma\sqrt{2\pi}$, respectively.
@@ C.2.2 Solution
Finally, using these relations in Eqs. (60, 61), we obtain
@@ C.3.1 Pair potential in Fourier space
The pair correlation function $h_{11}(r)$ is plotted in Fig. 3 for different values of the temperature $T_{2}$: a good agreement is obtained, both at equilibrium and out of equilibrium.
Figure 3: Pair correlation function $h(r)=g(r)-1$ of particles of the species 1, with $\rho=10$, $T_{1}=1$, for different temperatures $T_{2}$.
@@ D.3.2 Order $\epsilon^{2}$
Using the decomposition (95), we have to determine the integral of $w^{(2)}_{123}$, $w^{(2)}_{231}$ and $w^{(2)}_{312}$ over $\bm{x}_{3}$:
$\int w^{(2)}_{123}\text{d}\bm{x}_{3}$ $=\int\omega^{(2)}_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (118)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}^{(2)}_{123}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{2}-\bm{x}_{1})+i\bm{k}^{\prime}\cdot(\bm{x}_{3}-\bm{x}_{1})}\text{d}\bm{x}_{3}$ (119)
$=\int\frac{\text{d}\bm{k}}{(2\pi)^{d}}\tilde{\omega}^{(2)}_{123}(\bm{k},0)e^{i\bm{k}\cdot(\bm{x}_{2}-\bm{x}_{1})}$ (120)
@@
$\int w^{(2)}_{231}\text{d}\bm{x}_{3}$ $=\int\omega^{(2)}_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (122)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}^{(2)}_{231}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{3}-\bm{x}_{2})+i\bm{k}^{\prime}\cdot(\bm{x}_{1}-\bm{x}_{2})}\text{d}\bm{x}_{3}$ (123)
$=\int\frac{\text{d}\bm{k}^{\prime}}{(2\pi)^{d}}\tilde{\omega}^{(2)}_{231}(0,\bm{k}^{\prime})e^{i\bm{k}^{\prime}\cdot(\bm{x}_{1}-\bm{x}_{2})}$ (124)
@@
$\int w^{(2)}_{312}\text{d}\bm{x}_{3}$ $=\int\omega^{(2)}_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (126)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{\omega}^{(2)}_{312}(\bm{k},\bm{k}^{\prime})\int e^{i\bm{k}\cdot(\bm{x}_{1}-\bm{x}_{3})+i\bm{k}^{\prime}\cdot(\bm{x}_{2}-\bm{x}_{3})}\text{d}\bm{x}_{3}$ (127)
$=\int\frac{\text{d}\bm{k}}{(2\pi)^{d}}\tilde{\omega}^{(2)}_{312}(\bm{k},-\bm{k})e^{i\bm{k}\cdot(\bm{x}_{1}-\bm{x}_{2})}$ (128)
@@
$\hat{g}_{2,2}=-\int\phi^{(2)}\text{d}\bm{x}_{3}=-\tau_{312}\int w^{(2)}_{312}\text{d}\bm{x}_{3}=\frac{\tau_{312}}{2T_{12}}v_{31}*v_{32}(\bm{x}_{1}-\bm{x}_{2}).$ (131)
@@ D.3.3 Order $\epsilon^{3}$
Using the decomposition (95), we have to determine the integral of the six terms $v_{13}w_{123}^{(2)}$, etc. over $\bm{x}_{3}$. They are (we use $\bm{x}=\bm{x}_{1}-\bm{x}_{2}$)
$\int v_{13}w^{(2)}_{123}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega^{(2)}_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (135)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{123}(\bm{k},\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (136)
@@
$\int v_{23}w^{(2)}_{123}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega^{(2)}_{123}(\bm{x}_{2}-\bm{x}_{1},\bm{x}_{3}-\bm{x}_{1})\text{d}\bm{x}_{3}$ (137)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k})\tilde{\omega}^{(2)}_{123}(\bm{k}^{\prime},\bm{k})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (138)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{123}(\bm{k}-\bm{k}^{\prime},\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (139)
@@
$\int v_{13}w^{(2)}_{231}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega^{(2)}_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (140)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k})\tilde{\omega}^{(2)}_{231}(\bm{k},\bm{k}^{\prime})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (141)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{231}(\bm{k}^{\prime},\bm{k}-\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (142)
@@
$\int v_{23}w^{(2)}_{231}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega^{(2)}_{231}(\bm{x}_{3}-\bm{x}_{2},\bm{x}_{1}-\bm{x}_{2})\text{d}\bm{x}_{3}$ (143)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{231}(\bm{k}^{\prime},\bm{k})e^{i\bm{k}\cdot\bm{x}}.$ (144)
@@
$\int v_{13}w^{(2)}_{312}\text{d}\bm{x}_{3}$ $=\int v_{13}(\bm{x}_{1}-\bm{x}_{3})\omega^{(2)}_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (145)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k})\tilde{\omega}^{(2)}_{312}(\bm{k}^{\prime},-\bm{k}-\bm{k}^{\prime})e^{i(\bm{k}+\bm{k}^{\prime})\cdot\bm{x}}$ (146)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{13}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{312}(-\bm{k}-\bm{k}^{\prime},\bm{k})e^{i\bm{k}\cdot\bm{x}}.$ (147)
@@
$\int v_{23}w^{(2)}_{312}\text{d}\bm{x}_{3}$ $=\int v_{23}(\bm{x}_{2}-\bm{x}_{3})\omega^{(2)}_{312}(\bm{x}_{1}-\bm{x}_{3},\bm{x}_{2}-\bm{x}_{3})\text{d}\bm{x}_{3}$ (148)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k})\tilde{\omega}^{(2)}_{312}(\bm{k}^{\prime},-\bm{k}-\bm{k}^{\prime})e^{i\bm{k}^{\prime}\cdot\bm{x}}$ (149)
$=\int\frac{\text{d}\bm{k}\text{d}\bm{k}^{\prime}}{(2\pi)^{2d}}\tilde{v}_{23}(\bm{k}^{\prime})\tilde{\omega}^{(2)}_{312}(\bm{k},-\bm{k}-\bm{k}^{\prime})e^{i\bm{k}\cdot\bm{x}}.$ (150)
@@ D.4 Scaling relations and numerical integration
D.4.1 Scaling form
All the terms that involve the many-body potential $\phi$ at order $\epsilon^{3}$ consist of a double integral in Fourier space, three interaction potentials and a scale-invariant combination of the wavevectors (which scales as $k^{0}$). While the prefactor is difficult to obtain analytically, this form suggests an algebraic decay as $r^{-2d}$, where $d$ is the spatial dimension, and a prefactor scaling as $\tilde{V}(0)^{3}$, where $\tilde{V}(0)=\int V(\bm{x})\text{d}\bm{x}$. This is confirmed by the numerical integration for the Harmonic and Gaussian potentials presented in Sec. D.4.2 (Fig. 4).
@@ D.4.1 Scaling form
The dependence in the potential should be accompagned with a dependence in the temperatures, as only the ratio matters, as can be seen from the definition of the Liouville operator (Eq. (78)). If the particles 1 and 2 are of type $A$ and the others are of type $B$, these arguments leads to an asymptotic form
$\hat{g}(r)\underset{r\to\infty}{\sim}\left[\frac{\tilde{V}(0)}{T_{A}}\right]^{3}F\left(\frac{T_{B}}{T_{A}}\right)r^{-2d}.$ (155)
As the correction originates from temperature differences, the temperature dependence should satisfy
$F(\theta)\underset{\theta\to 1}{\sim}\theta-1.$ (156)
Moreover counting the occurence of the temperatures $T_{A}$ and $T_{B}$ in the different terms entering the correction, we find that for large temperature differences,
$F(\theta)\underset{\theta\to\infty}{\sim}\theta^{-2}.$ (157)